OpenStudy (anonymous):

Find f(-2) if f(x) = 2x2 + kx + 2 and f(1) = 3.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

replace x by 1 and solve \[f(1)=2+k+2=3\] for k

OpenStudy (anonymous):

k=-1?

OpenStudy (anonymous):

no it is wrong

OpenStudy (anonymous):

We find what k is by plugging 1nfor x into the original function and setting the result equal to 3: 2*1^2 + k*1 + 2 = 3, so k = -1, and the function is\[f(x) = 2x^{2} - x + 2.\]We evaluate this function at x = -2: f(-2) = 8.

OpenStudy (anonymous):

so what is its' result?

OpenStudy (anonymous):

I glitched. f(-2) = 12.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Did you get k = -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

looks like if you solve larger less than twice the smaller. make x the larger then we get \[2+k+2=3\] you get \[k=-1\] yes? so now you have \[f(x)=2x^2-x+2\] and if you want \[f(-2)\] you compute \[f(-2)=2(-2)^2-(-2)+2\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

You use the information to find the value of f(-2).

OpenStudy (anonymous):

ok I understand now.Thanks

OpenStudy (anonymous):

:^)